OpenStudy (anonymous):

solve cot^-1(cot(-pi/8)) without a calculator??

OpenStudy (wwe123):

Cot( x/2)=cos(x/2) / sin(x/2) = sqr(1+cosx)/sqr(1-cosx) =sinx/(1-cosx) So cot (pi/8)= sin (pi/4) / [1-cos(pi/4)] =(sqr2)/2 / [1-(sqr2)/2] =sqr2/(2-sqr2) Sqr2(2+sqr2)/2 = (2sqr2+2)/2 = sqr2+1

OpenStudy (cggurumanjunath):